OpenStudy (luigi0210):

Find the Limit:

OpenStudy (luigi0210):

\[\lim_{x \rightarrow -2} \frac{x+2}{x^3+8}\]

zepdrix (zepdrix):

`Sum of Cubes Formula`:\[\Large a^3+b^3 \qquad=\qquad (a+b)(a^2-ab+b^2)\]That will allow us to change our denominator a bit :)

zepdrix (zepdrix):

\[\Large x^3+8\qquad=\qquad x^3+2^3\qquad=\qquad (x+2)(x^2-2x+2^2)\]Understand how that works? :o

OpenStudy (luigi0210):

Haha, yes @zepdrix :) Couldn't we also use L'hoital's rule?

zepdrix (zepdrix):

Since we're getting the indeterminate form 0/0, yes L'Hop would work just fine! :)

OpenStudy (anonymous):

\[\lim_{X \rightarrow -2}\frac{x+2}{x^3+8}=\lim_{X \rightarrow -2}\frac{x+2}{x^3+2^3}\] \[=\lim_{X \rightarrow -2}\frac{x+2}{(x+2)(x^2-2x+2^2}=\lim_{X \rightarrow -2}\frac{x+2}{(x+2)(x^2-2x+4}\] \[=\lim_{X \rightarrow -2}\frac{1}{(x^2-2x+4}\] Applying limit \[=\frac{1}{(-2)^2-2(-2)+4} =\frac{1}{4+4+4} =\frac{1}{12} \]

OpenStudy (luigi0210):

Thanks guys :) @zepdrix @dpasingh